Ch 2 HW Compute the present values of the following annuities frst assuming that payments are made o Saved Help Save & Exit Submit Check my work Problem 2-27 (LG 2-9) Compute the present values of the following annuities first assuming that payments are made on the last day of the period and then assuming payments are made on the first day of the period: (Do not round intermediate calculations. Round your answers to 2 decimal places. (e.g., 32.16) Present Value Present Value Payment Years Interest Rate s 778.09 Years (Annuall Payment made on (Payment made on first day of period) last day of period) 81 141% 14 8,968 26 21.32.93 24 70,712 54 32 Next > Prev5 of 15

Ch 2 HW Saved Help Save & Exit Submit Check my work 14 Problem 2-25 (LG 2-9) Calculate the future value of the following annuity streams: points a. $5,000 received each year for 7 years on the last day of each year if your investments pay 7 percent compounded annually. b. $5,000 received each quarter for 7 years on the last day of each quarter if your investments pay 7 percent compounded quarterly c. $5.000 received each year for 7 years on the first day of each year if your investments pay 7 percent compounded annually. d. $5,000 received each quarter for 7 years on the first day of each quarter if your investments pay 7 percent compounded eBook Print References quarterly (For all requirements, do not round intermediate calculations. Round your answers to 2 decimal places. (e.g. 32.16) a. Future value b. Future value c. Future value d. Future value < Prev 14 of 15 ill Next >
Answer #1

In general, Present Value formula: PV = A/(1+r) + A/(1+r)^2 + ...+ A/(1+r)^n ; A =annuity; r = yearly interest rate, n = number of years.

Assuming payments made at the end of the period (EOP)

  • Annuity, A = 778.09; for years = 8 at rate = 14%
    • PV =778.09/(1+0.14) + 778.09/(1+0.14)^2 + .... + 778.09/(1+0.14)^8 = $3,609.45
  • Annuity, A = 8,968.26; for years = 14 at rate = 7%
    • PV = 8,968.26/(1+0.07) + 8,968.26/(1+0.07)^2 + .... + 8,968.26/(1+0.07)^14 = $78,431.63
  • Annuity, A = 21,322.93; for years = 24 at rate = 5%
    • PV = 21,322.93/(1+0.05) + 21,322.93/(1+0.05)^2 + .... + 21,322.93/(1+0.05)^24 = $294,227.47
  • Annuity, A = 70,712.54; for years = 5 at rate = 32%
    • PV = 70,712.54/(1+0.32) + 70,712.54/(1+0.32)^2 + .... + 70,712.54/(1+0.32)^5 = $165,835.41

Assuming payments made at the beginning of the period (BOP)

  • Annuity, A = 778.09; for years = 8 at rate = 14%
    • PV =778.09 + 778.09/(1+0.14) + 778.09/(1+0.14)^2 + .... + 778.09/(1+0.14)^7 = $4,114.78
  • Annuity, A = 8,968.26; for years = 14 at rate = 7%
    • PV = 8,968.26 + 8,968.26/(1+0.07) + 8,968.26/(1+0.07)^2 + .... + 8,968.26/(1+0.07)^13 = $83,921.84
  • Annuity, A = 21,322.93; for years = 24 at rate = 5%
    • PV =21,322.93 + 21,322.93/(1+0.05) + 21,322.93/(1+0.05)^2 + .... + 21,322.93/(1+0.05)^23 = $308,938.85
  • Annuity, A = 70,712.54; for years = 5 at rate = 32%
    • PV = 70,712.54 + 70,712.54/(1+0.32) + 70,712.54/(1+0.32)^2 + .... + 70,712.54/(1+0.32)^4 = $218,902.74

Summary:

Annuity Years rate PV calc at EOP PV calc at BOP
778.09 8 14% $3,609.45 $4,114.78
8968.26 14 7% $78,431.63 $83,921.84
21,322.93 24 5% $294,227.47 $308,938.85
70,712.54 5 32% $165,835.41 $218,902.74

Future Value:

In general, Future Value formula = A(1+r) + A(1+r)^2 + ... + A(1+r)^n

[EOP = End of period, BOP = Beginning of period]

a) A=5000 invested at the EOP, n=7 years, rate = 7%, compounded annually

FV = 5000(1+0.07)^6 + 5000(1+0.07)^5 + ...+ 5000(1+0.07)^1 + 5000 = $43,270.11

b) A=5000 invested at the EOP, n=7 years, rate = 7%, compounded quarterly

number of compounding periods = 7 * 4 = 28; compounding rate = 7%/4 =  1.75%

FV = 5000(1+0.0175)^27 + 5000(1+0.0175)^26 + ...+ 5000(1+0.0175)^1 + 5000 = $178,688.40

c) A=5000 invested at the BOP, n=7 years, rate = 7%, compounded annually

FV = 5000(1+0.07)^7 + 5000(1+0.07)^6 + ...+ 5000(1+0.07)^1 = $46,299.01

d) A=5000 invested at the BOP, n=7 years, rate = 7%, compounded quarterly

number of compounding periods = 7 * 4 = 28; compounding rate = 7%/4 =  1.75%

FV = 5000(1+0.0175)^28 + 5000(1+0.0175)^27 + ...+ 5000(1+0.0175)^1 = $178,691.02

Summary:

Annuity investment period annual rate Compounding Invested at Future Value
5000 7 7% annual EOP $43,270.11
5000 7 7% quarterly EOP $178,689.40
5000 7 7% annual BOP $46,299.01
5000 7 7% quarterly BOP $178,691.02